140813 LML Observations

Isaac arrived at LML at 9:50 am. The doors to the collector had been deployed but no sample was present. The collector was then disassembled and the collector components were cleaned with 409 and DI water. The collector body was then cleaned with 409 and rinsed the DI water as well. The collector was then reassembled and reset for the next fog event.

Leave a Reply

Your email address will not be published. Required fields are marked *